Maricopa's Success scholarship fund receives a gift of $ 210000. The money is invested in stocks, bonds, and CDs. CDs pay 4.5 % interest, bonds pay 4.6 % interest, and stocks pay 7.2 % interest. Maricopa Success invests $ 20000 more in bonds than in CDs. If the annual income from the investments is $ 11420 , how much was invested in each account?
Answer by VFBundy(438)

CDs:
Principal = p
Rate = 0.045
Interest = 0.045p

Bonds:
Principal = p + 20000
Rate = 0.046
Interest = 0.046(p + 20000) = 0.046p + 920

Stocks:
Principal = 210000 - p - (p + 20000) = 190000 - 2p
Rate = 0.072
Interest = 0.072(190000 - 2p) = 13680 - 0.144p

Total interest:
(0.045p) + (0.046p + 920) + (13680 - 0.144p) = 11420

Simplify:

-0.053p + 14600 = 11420

-0.053p = -3180

p = 60000

From earlier:

CDs:
Principal = p = 60000

Bonds:
Principal = p + 20000 = 60000 + 20000 = 80000

Stocks:
Principal = 190000 - 2p = 190000 - 120000 = 70000